# Edge density and minimum degree thresholds for $H$-free graphs with unbounded chromatic number

### v1

@@ 1 Introduction
The bounds presented in these theorems are sharp, as demonstrated by explicit constructions based on Erdős graphs and blowups of Borsuk–Hajnal graphs (see Figure 2).
@@ 2.1 Extremal graphs
Let $\mathrm{BH}^{-}$ be the graph on vertex set $U^{\prime}\cup X$ obtained from $\mathrm{BH}(n,k,\ell,\alpha)$ by removing the vertex set $W$. We define two families of graphs, $\mathrm{BH}_{r,\delta}^{*}$ and $\mathrm{BH}_{r,\delta}^{**}$, which are obtained from $BH^{-}$ by adding vertex sets $Y=Y_{1}\cup\cdots\cup Y_{r-3}$ and $Z$. The edges are added as follows:
@@
  • For $\mathrm{BH}_{r,\delta}^{*}:|Y_{i}|=\frac{1-\delta}{2\delta-2(r-3)(1-\delta)}|X|$ and $|Z|=\frac{1-2\delta+(r-3)(1-\delta)}{2\delta-2(r-3)(1-\delta)}|X|$.
  • For $\mathrm{BH}_{r,\delta}^{**}:|Y_{i}|=\frac{6\delta-1}{2(\delta(5r-14)-(r-3)(6\delta-1))}|X|$ and $|Z|=\frac{(1-2\delta)(5r-14)+(r-3)(6\delta-1)}{2(\delta(5r-14)-(r-3)(6\delta-1))}|X|$.
@@ 2.2 The Regularity lemma
Let $G$ be a graph on $n$ vertices. For given $\varepsilon>0$ and $d\in(0,1)$, let $R$ be the corresponding ($\varepsilon,d$)-reduced graph. Then $e(G)\leq\left(e(R)+o(1)\right)\frac{n^{2}}{k^{2}}$.

### v2

@@ 1 Introduction
Moreover, $\delta_{\chi}(H)\neq\frac{r-2}{r-1}$ if and only if $H$ has a forest in its decomposition family, and $\delta_{\chi}(H)=\frac{r-3}{r-2}$ if and only if $H$ is $r$-near-acyclic. (For the precise definitions of the decomposition family and the concept of being $r$-near-acyclic, see Section 2.) For further recent progress, see bourneuf2025denseneighborhoodlemmaapplications ; huang2025interpolating ; liu2024beyond ; xue2025directed .
@@
Remark.
The bound in Theorem 1.1 is asymptotically tight; explicit constructions based on blowups of Borsuk–Hajnal graphs matching these densities are provided in Section 2.1 (see Figure 2 (a) and (b)). Theorem 1.1 can be stated as a sufficient condition for bounded chromatic number in terms of a “density-deficit” trade-off. Let $\rho=e(G)/n^{2}$ be the edge density. Theorem 1.1 implies that if
$h(\rho,\delta):=\rho+A(\delta-B)^{2}>C,$
then $\chi(G)$ is bounded. Notably, this trade-off is linear in the global density $\rho$ but quadratic in the local parameter $\delta$. The parameters $A,B,C$ depend on the range of $\delta$:
  • Upper Regime ($\frac{5r-13}{5r-8}<\delta\leq\frac{2r-5}{2r-3}$): In this range, the parameters are given by:
    $A=\frac{5r^{2}-17r+14}{2},\quad B=\frac{5r^{2}-22r+23}{5r^{2}-17r+14},\quad C=\frac{5r^{2}-22r+25}{2(5r^{2}-17r+14)}.$
  • Lower Regime ($\frac{2r-5}{2r-2}\leq\delta\leq\frac{5r-13}{5r-8}$): In this range, the parameters are given by:
    $A=\frac{2r-2}{5r-14},\quad B=\frac{2r-5}{2r-2},\quad C=\frac{r-2}{2r-2}.$
@@
The bound presented in Theorem 1.2 are sharp, as demonstrated by explicit constructions based on Erdős graphs (see Figure 2 (c)).
@@ 2.1 Extremal graphs
Let $\mathrm{BH}^{-}$ be the graph on vertex set $U^{\prime}\cup X$ obtained from $\mathrm{BH}(n,k,\ell,\alpha)$ by removing the vertex set $W$. We define two families of graphs, $\mathrm{BH}_{r,\delta}^{\star}$ and $\mathrm{BH}_{r,\delta}^{\star\star}$, which are obtained from $BH^{-}$ by adding vertex sets $Y=Y_{1}\cup\cdots\cup Y_{r-3}$ and $Z$. The edges are added as follows:
@@
  • For $\mathrm{BH}_{r,\delta}^{\star}:|Y_{i}|=\frac{1-\delta}{2\delta-2(r-3)(1-\delta)}|X|$ and $|Z|=\frac{1-2\delta+(r-3)(1-\delta)}{2\delta-2(r-3)(1-\delta)}|X|$.
  • For $\mathrm{BH}_{r,\delta}^{\star\star}:|Y_{i}|=\frac{6\delta-1}{2(\delta(5r-14)-(r-3)(6\delta-1))}|X|$ and $|Z|=\frac{(1-2\delta)(5r-14)+(r-3)(6\delta-1)}{2(\delta(5r-14)-(r-3)(6\delta-1))}|X|$.
@@
Remark.
Let $H$ be a graph with $\chi(H)=r$ that is not $r$-near-acyclic, which corresponds to the threshold $\delta_{\chi}(H)=\frac{2r-5}{2r-3}$. We note that our extremal constructions, $\mathrm{BH}^{\star}_{r,\delta}$ and $\mathrm{BH}^{\star\star}_{r,\delta}$, are derived from the $r$-Borsuk–Hajnal graph by adjusting the sizes of its independent sets. The underlying $r$-Borsuk–Hajnal graph was shown to be $H$-free in allen2013chromatic , providing the structural basis for our construction.
@@
Remark.
Let $H$ be a graph with $\chi(H)=r$. We verify that our construction $\mathrm{EG}_{r,\delta}$ is indeed $H$-free. Since the Erdős graph component is chosen to have a large girth, any subgraph induced by $|H|$ vertices in $V(E)\cup X\cup Z$ is $2$-colorable. Hence, any subgraph induced by $|H|$ vertices in the entire graph $\mathrm{EG}_{r,\delta}$ is (r1)(r-1)-colorable, which implies that $\mathrm{EG}_{r,\delta}$ is $H$-free.
@@ 2.2 The Regularity lemma
For $\varepsilon>0$ and $d\in(0,1)$, let $G$ be a graph on $n$ vertices with an ($\varepsilon,d$)-reduced graph $R$ on $k$ vertices, where $1/k\leq 2\varepsilon$. Then $e(G)\leq e(R)\frac{n^{2}}{k^{2}}+\left(\frac{d}{2}+2\varepsilon\right)n^{2}$.
Proof.
The edges of $G$ can be categorized into the following four types:
  • Edges within clusters: The number of edges with both endpoints in $V_{i}$ is at most $k\binom{n/k}{2}$;
  • Edges in irregular pairs: The number of edges between distinct clusters $V_{i}$ and $V_{j}$ such that $(V_{i},V_{j})$ is not $\varepsilon$-regular is at most $\varepsilon k^{2}\frac{n^{2}}{k^{2}}$;
  • Edges in sparse regular pairs: The number of edges between distinct clusters $V_{i}$ and $V_{j}$ where $(V_{i},V_{j})$ is $\varepsilon$-regular but $d(V_{i},V_{j})<d$ is at most $d\binom{k}{2}\frac{n^{2}}{k^{2}}$;
  • Edges in dense regular pairs: The edges between distinct clusters $V_{i}$ and $V_{j}$ such that $(V_{i},V_{j})$ is $\varepsilon$-regular and $d(V_{i},V_{j})\geq d$ correspond to the edges of $R$. The number of such edges is at most $e(R)\frac{n^{2}}{k^{2}}$.
Summing the upper bounds for these four types yields the desired result. ∎